OpenStudy (anonymous):

Help, please! :) Choose the correct simplification of the expression the quantity a to the 3rd power over b to the 7th power all raised to the 2nd power. a to the 5th power over b to the 9th power a over b to the 14th power a to the 6th power over b to the 14th power a6b14 10. [7.04] Choose the correct simplification of the expression 6 times b all over a to the power of negative 9. a to the 9th power times b all over 6 6 times b all over a to the 9th power Already simplified 6a9b

OpenStudy (wmj259):

\[(\frac{ a ^{3}\ }{ b ^{7} } )^2\]

OpenStudy (wmj259):

Like that?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

@johnweldon1993 could you help? :)

OpenStudy (anonymous):

\[\frac{a^{3*2}}{b^{7*2}}\]

OpenStudy (anonymous):

sup?

OpenStudy (anonymous):

or you can think of it like this \[\frac{a^3}{b^7}*\frac{a^3}{b^7}\]